OpenStudy (anonymous):

is √2x is equivalent to √x2

OpenStudy (anonymous):

is that x(2) or is it \(x^2\)

OpenStudy (anonymous):

square root of 2x and square root of x2

OpenStudy (anonymous):

Is that x2 supposed to be x squared? or is it supposed to be x times 2?

myininaya (myininaya):

\[\sqrt{2x}=\sqrt{x2},\sqrt{2x} \neq \sqrt{x^2}\]

OpenStudy (anonymous):

it's not x times two, or x squared. it's just x2? lol

myininaya (myininaya):

x2 is x times 2

OpenStudy (anonymous):

unless its \(x_2\)

myininaya (myininaya):

lol

myininaya (myininaya):

what is x_2

OpenStudy (anonymous):

oh,lol. so true or false?

OpenStudy (anonymous):

whatever you want it to be.

myininaya (myininaya):

x_2=polpak?

OpenStudy (anonymous):

if x2 is x times 2 then it's true. If x2 is \(x_2\) then it's anyone's guess. If x2 is \(x^2\) then it's false.

myininaya (myininaya):

and i have what he said above with the exception of the x_2

myininaya (myininaya):

because for some reason i believe it is not that

myininaya (myininaya):

so that 2 after the x does it looked raised on your paper?

myininaya (myininaya):

if it looks raised then its an exponent if it is the same font size and same format then it is x times 2

OpenStudy (anonymous):

it's not raised at all. but thanks guys for helping, i already finished :)